OpenStudy (anonymous):

please help. Find the area of the indicated region Between y = x^2 and y = −1 for x in [−1, 1]

OpenStudy (anonymous):

did you draw the region?

OpenStudy (anonymous):

I graphed it

OpenStudy (anonymous):

or is the region already given to you

OpenStudy (anonymous):

it'd just be \[A=\int_{x=-1}^1(1-x^2)dx\]

OpenStudy (anonymous):

wait.. is that \[\Large y=-1\qquad{\rm OR}\qquad y=1 \]

OpenStudy (anonymous):

y= -1

OpenStudy (anonymous):

then, the other graph must be\[\Large y=-x^2 \qquad{\rm NOT}\quad y=x^2\]

OpenStudy (anonymous):

it says y = x^2

OpenStudy (anonymous):

post the picture of the graph

OpenStudy (anonymous):

but if that is the case, then \[A=\int_{x=-1}^1(x^2+1)dx\]

OpenStudy (anonymous):

 the graph looks like this

OpenStudy (anonymous):

ok. then it is the second integration I just wrote